# Reinforcement Learning for Control of Evolutionary and Ecological Processes

### v2

@@ DISCUSSION
Though a discussion of the practical learning procedure itself is beyond the scope of this study given it will likely be a contextual decision based on the desired application of directed evolution (clinical, industrial, scientific, etc.), our theoretical analysis stands on its own as a first-of-its kind approach for addressing one of the central questions of applied evolutionary theory. The interoperability of Computational duality makes it particularly suitable for the monitoring and course-correction of evolutionary processes, given the multicules corresponding to an intercellular compartment provide a ready readout for the state of the system provided there are adequately accurate inference methods for translating back and forth with experimental data. We expect such inference methods to be very feasible.
@@
We see a number of promising directions for future work, including 1) extending the results to nonlinear transforms $c=\mathbf{B}(x)$ rather than the linear one we assume, 2) analyzing the potential impacts of the difference between the timescale of the respective dynamics of $x$ and $u$, 3) tightening the bounds on quadrature estimate by accounting for sparcity in parameter matrices, and of course 5) applying our approach or variants of it to experimental problems.
@@
In summary, this study represents three key advancements. First, a novel mathematical formulation of Computational duality for biophysical dynamics which seamlessly links ecology and evolution into a single unified process – just as is of course the reality in nature. Second, Computational duality’s universality and ability to more richly interface with the mechanistic properties of biosystems has the potential to strengthen the link between theory and experiment in further highly novel ways. Third, by establishing for the first time the utility of tools from modern artificial intelligence theory cross-applied to the mathematical study of evolution and the biophysical questions therein, we find that in addition to the established relationship of neuroscience and AI this offers a new explicit link between biophysical processes and that of artificial intelligence. All three represent interesting directions for future work of a highly interdisciplinary nature.
@@
Because we sought to provide a model that could interact directly with experiments, certain decisions were made which increase the conceptual and computational complexity of the problem. For example, the mathematical model itself is notably more complex than those typically used in evolutionary dynamics. However, each of the modeling decisions was carefully made so that as much prior experimental data as possible could be incorporated into the model before any training might begin. Our goal was to strike a balance between the capacity for integration with physiological experiments – where simpler models have often struggled – and the need to retain a model tractable for algorithmic analysis so that rigorous statements could be made about an interpretable model. Proving bounds for PSRL is straightforward for the coarse-grained models typical in evolutionary dynamics, so in a sense that result is incorporated here as a limit case of our work.
Because the values of $\varepsilon_{O}$ and $\varepsilon_{\text{min}}$ are difficult to estimate in the absence of experimental data, this represents a key aspect which we do not presently know of any means of analyzing mathematically a priori. Methods such as those described in the above discussion of Reduced-Order Models may offer useful tools, but will require substantial calculations in order to prove bounds on $\varepsilon_{O}$. By contrast, $\varepsilon_{\text{min}}$ will almost certainly need to be determined experimentally for each evolving system which one might wish to study. Because of this, it is entirely possible and perhaps inevitable that our approach and derivatives of it will be useful for certain evolving systems, but not prove effective for others.
@@ METHODS
Evolutionary game theoretic models are necessarily nonlinear dynamical systems, and in practice at present it is generally only possible to take estimates of their states occasionally due to the need for genetic sequencing or molecular testing of some sort. Thus, any effort to direct evolution will require the application of digital control.33, 34 In reality, $\mathbf{F}$ is a phenomenological representation of some underlying nonlinear system describing the dynamics of biochemistry and signalling processes within and between biological cells and organisms. Recent improvements in biological data-gathering technology 35, 36 enable the inference of a linear approximation $\mathbf{G}x$ of an underlying nonlinear system for an $n_{0}$-dimensional vector $x$ of molecular levels within cells. It should be noted that $\mathbf{G}$ will be a sparse matrix, as any given molecule will typically only interact meaningfully with a small number of other molecules. Exceptions absolutely exist, such as the central importance of glucose [^7] A sugar which serves as the body’s main source of energy system-wide. and ATP [^8] A high-energy molecule synthesized from a variety of energy sources, which is used to fuel a wide array of cellular processes., but by and large $\mathbf{G}$ will be quite sparse. [^9] That is, most of the elements of the matrix will be zero.
The complexity of biological systems presents substantial challenges, but this complexity also produces enormous opportunities if methods can be developed for deliberately and precisely controlling them.37, 38, 36 Given there are modeling principles which it is believed should transfer over to complex, data-driven, real-world cases, Model Predictive Control appears to be the most viable option for solving this complex MIMO [^10] Multiple Input, Multiple Output control problem.39, 40
@@
Specific $\vec{\omega}$ and $\vec{\mu}$ are determined by the goals in some problem of interest. In this study these goals are set by the humans aiming to control some eco-evolutionary process, but extensions to this model could define it based on an ecology itself to study natural evolution. [^12] For instance, one could set $u(k)$ to zero, use a different ecology function than we do here, and have $R(x(k))$ be simply maximization over replication rates. In other words, survival of the fittest.
There are a few options in terms of how to model a system where data is available in the form of values for $c$, while the dynamics themselves are in truth defined in terms of $x$. One option is to assume a given estimate of $c$ maps onto an estimated $x$ (and vice versa), and have the model be a hybrid of dynamics over $c$ and $x$. For the purposes of this study, the approach will be to assume a pseudo-invertable [^13] See Section 1 for a discussion of the validity of the assumption of pseudo-invertability. linear map $\mathbf{B}x=c$ which is time-invariant. Incorporating noise will be relevant later, so we will use the notation $\mathbf{\tilde{B}}=\mathbf{B}x+\xi$ with $\xi$ being a vector of zero-mean noise. For the time being, however, we will assume perfect ”estimates” and ignore the $\xi$ component of $\mathbf{\tilde{B}}$. Further, we define a vector of inputs to the ecology $u$, and a matrix $\mathbf{Z}$ mapping inputs to multicules.
However, it is not in fact required that we have any knowledge of $\mathbf{B}$, $\mathbf{G}$, or any other parameter matrices a priori. This is where the application of reinforcement learning will become central to later sections of the study. For now, we must first define a model of the eco-evolutionary process. The rest of this section will be a discussion of the model and how we can convert it into a Partially Observable Markov Decision Process (POMDP) so as to enable the use of the extensive research in control systems engineering and the theory of artificial intelligence.
@@
As mentioned in the introduction, the complexity of eco-evolutionary processes can be enormous. Tracking the positions of even a negligible fraction of molecules involved in such processes is not only infeasible but almost certainly unproductive. However, there is value in being able to capture key features of spatial dynamics as well as more direct biochemical and evolutionary ones. One simplified model which offers a good balance between the two is to analyze the spatial component of a system as a set of discrete partitions.
We can imagine attaching to any molecule a hypothetical string of bits to serve as an address of sorts. One value for the bit string could say that it is within a biological cell, another that it is outside the cell. If more fidelity is desired, one could denote intranuclear, [^14] That is, within the nucleus of a cell. cytosolic, [^15] Within the region between a cell’s nucleus and its outer boundary. or extrecellular. Alternatively (or going further), one could create separate label string values for any given partition of interest, such as one value for the shared space between a group of cells and then additionally one value each for every type of cell as described in LABEL:section_background_defining_egt.
@@
Computational duality allows us to capture richer spatial aspects of eco-evolutionary dynamics in the molecular reformulation of EGT discussed above, as we can account for molecule placement simply by creating multiple entries in $x$ which correspond to the same physical molecular structure across as many partitions as relevant for a given problem. For $q$ cell types of interest, one extracellular compartment, and $n_{0}$ distinct molecular structures of interest, we can have up to $n_{\text{max}}=n_{0}(q+1)$ multicules. It is a set of $n\leq n_{\text{max}}$ multicules which we consider here.
Computational duality is a particularly interesting proposition in that is is capable of converting essentially any collection of disparate types of biological processes into a single coherent process type. Abstracting small-molecule structural information and position as part of a collective state can, in principle, be extended to discretizations of other properties such as conformational states of proteins, methylation states of DNA. Versions of this concept appear in systems theory in various contexts, and biophysics allows all properties to be represented as abstract analogs to molecular structures, and all processes as abstract analogs to chemical reactions. The general nature of this view is the reason for the term multicule rather than particule. This combined state space is of course computationally intractably large if one wished to include the entirety of most given biological processes, but reduced state spaces can always be chosen for the variables most relevant to a given scientific question. Perhaps even more interesting is that from a mathematical perspective generalized multicules provide a single, unified framework for all of biology.
@@
  • Because this study focused on the mathematical aspects of an applied theory problem in mathematical biology, no code or experimental data were used.
@@ ACKNOWLEDGMENTS
This work was funded by a research grant from the Stanford University Physician-Scientist Training Program. We would like to thank Mykel Kochenderfer, Arec Jamgochian, and Benjamin van Roy for guidance in literature review and recommendations of relevant papers in the fields of POMDPs and RL, and the former two for helpful feedback on conceptualization in the early stages of the project. Additionally, we thank the late Christopher T. Walsh for valuable discussions of underlying concepts in the early stages of the project and for his mentorship. He will be missed.
@@
DECLARATION OF GENERATIVE AI AND AI-ASSISTED TECHNOLOGIES
During the preparation of this work, the authors made no use of Generative AI or AI-assisted technology.
@@
2: Model-Predictive Control of the Eco-evolutionary Game. This section describes the Bellman operator and Bellman function, particularly as relates to this paper.

### v3

@@ DISCUSSION
Thus far there has been relatively little work applying contemporary advancements in control theory and related disciplines to questions of evolutionary systems. There is a recent discussion of the use of dynamic programming for the optimization of cancer therapy,34 however, we have been unable to find any such work beyond that paper. The massive advances in the mathematical study of control systems and artificial intelligence from recent decades have not yet crossed over into the study of evolutionary dynamics and its application to cancer biology. This represents a fertile ground for the use of new tools, the answering of previously unstudied questions, and the framing entirely new questions.
@@
In this paper we expand the theoretical methods available for evolutionary dynamics via integrating contemporary research in control theory and the theory of artificial intelligence. Though a discussion of the experimental component of a practical learning procedure is beyond the scope of this study given it will likely be a contextual decision based on the desired application of directed evolution (clinical, industrial, scientific, etc.), our theoretical analysis stands on its own as a first-of-its kind approach for addressing one of the central questions of applied evolutionary theory. The interoperability of Computational duality makes it particularly suitable for the monitoring and course-correction of evolutionary processes, given the multicules corresponding to an intercellular compartment offer a ready readout for the state of the system provided there are adequately accurate inference methods for translating back and forth with experimental data. Cell sorting and whole-cell mass spectrometry likewise offer a means of reading out the state of intracellular multicules at least below the resolution level of organelles. We expect such inference methods to be very feasible.
@@
We see a number of promising directions for future work, including 1) extending the results to nonlinear transforms $c=\mathbf{B}(x)$ rather than the linear one we assume, 2) analyzing the potential impacts of the difference between the timescale of the respective dynamics of $x$ and $u$, 3) tightening the bounds on quadrature estimate by accounting for empirical features of parameter matrices, and of course 4) applying our approach or variants thereof to experimental problems.
The use of reinforcement learning means that there are minimal requirements of a-priori knowledge of a system. Beyond specifying definitions of multicules and input compounds, we could readily leave all parameter matrices unknown at the start of the problem. This will of course increase the number of trials needed to learn a model when compared with learning given partial prior knowledge of parameters, but the worst-case bounds described here are valid for the cases of any amount of prior knowledge, including the case of no such knowledge.
@@
In summary, this study represents three key advancements. First, a novel mathematical formulation of Computational duality for biophysical dynamics which seamlessly links ecology and evolution into a single unified process – just as is the reality in nature. Second, Computational duality’s universality and ability to more richly interface with the mechanistic properties of biosystems has the potential to strengthen the link between theory and experiment in further novel ways. Third, by establishing for the first time the utility of tools from modern artificial intelligence theory cross-applied to the mathematical theory of evolution and the biophysical questions therein, this offers a new explicit link between theoretical biophysics, ecology, evolutionary theory, and artificial intelligence. All three advancements represent interesting directions for future work of a highly interdisciplinary nature.
@@
Because we sought to provide a model that could interact directly with experiments, certain decisions were made which increase the conceptual and computational complexity of the problem. For example, the mathematical model itself is notably more complex than those typically used in evolutionary dynamics. However, each of the modeling decisions was carefully made so that as much prior experimental data as possible could be incorporated into the model before any training might begin. Our goal was to strike a balance between the capacity for integration with physiological experiments – where simpler models have often struggled – and the need to retain a model tractable for algorithmic analysis so that rigorous statements could be made about an interpretable model. Our results on PSRL learning bounds inherently translate to the coarse-grained models typical in evolutionary dynamics, so in a sense that result is incorporated here as the simplest limit case of our work.
@@ METHODS
Evolutionary game models are necessarily nonlinear dynamical systems, and in practice at present it is generally only possible to take estimates of their states occasionally due to the need for genetic sequencing or molecular testing of some sort. Thus, any effort to direct evolution will require the application of digital control.35, 36 In reality, $\mathbf{F}$ is a phenomenological representation of some underlying nonlinear system describing the dynamics of biochemistry and signaling processes within and between biological cells and organisms. Recent improvements in biological data-gathering technology 37, 38 enable the inference of a linear approximation $\mathbf{G}x$ of an underlying nonlinear system for an $n_{0}$-dimensional vector $x$ of molecular levels within cells. It should be noted that $\mathbf{G}$ will be a sparse matrix, as any given molecule will typically only interact meaningfully with a small number of other molecules. Exceptions absolutely exist, such as the central importance of glucose [^7] A sugar which serves as the body’s main source of energy system-wide. and ATP [^8] A high-energy molecule synthesized from a variety of energy sources, which is used to fuel a wide array of cellular processes., but by and large $\mathbf{G}$ will be quite sparse. [^9] That is, most of the elements of the matrix will be zero.
The complexity of biological systems presents substantial challenges, but this complexity also produces enormous opportunities if methods can be developed for deliberately and precisely controlling them.39, 40, 38 Given there are modeling principles which it is believed should transfer over to complex, data-driven, real-world cases, Model Predictive Control appears to be the most viable option for solving this complex MIMO [^10] Multiple Input, Multiple Output control problem.41, 42
@@
Specific $\vec{\omega}$ and $\vec{\mu}$ are determined by the goals in some problem of interest. In this study these goals are set by the humans aiming to control some evolutionary process, but extensions to this model could define it based on an ecology itself to study natural evolution. [^12] For instance, one could set $u(k)$ to zero, use a different ecology function than we do here, and have $R(x(k))$ be simply maximization over replication rates. In other words, survival of the fittest.
There are a few options in terms of how to model a system where data is available in the form of values for $c$, while the dynamics themselves are in truth defined in terms of $x$. One option is to assume a given estimate of $c$ maps onto an estimated $x$ (and vice versa), and have the model be a hybrid of dynamics over $c$ and $x$. For the purposes of this study, the approach will be to assume a pseudo-invertable [^13] See Supplemental Information Section 1 for a discussion of the validity of the assumption of pseudo-invertability. linear map $\mathbf{B}x=c$ which is time-invariant. Incorporating noise will be relevant later, so we will use the notation $\mathbf{\tilde{B}}=\mathbf{B}x+\xi$ with $\xi$ being a vector of zero-mean noise. For the time being, however, we will assume perfect ”estimates” and ignore the $\xi$ component of $\mathbf{\tilde{B}}$. Further, we define a vector of inputs to the ecology $u$, and a matrix $\mathbf{Z}$ mapping inputs to multicules.
However, it is not in fact required that we have any knowledge of $\mathbf{B}$, $\mathbf{G}$, or any other parameter matrices a priori. This is where the application of reinforcement learning will become central to later sections of the study. For now, we must first define a model of the evolutionary process. The rest of this section will be a discussion of the model and how we can convert it into a Partially Observable Markov Decision Process (POMDP) so as to enable the use of the extensive research in control systems engineering and the theory of artificial intelligence.
@@
As mentioned in the introduction, the complexity of evolutionary processes can be enormous. Tracking the positions of even a negligible fraction of molecules involved in such processes is not only infeasible but almost certainly unproductive. However, there is value in being able to capture key features of spatial dynamics as well as more direct biochemical and evolutionary ones. One simplified model which offers a good balance between the two is to analyze the spatial component of a system as a set of discrete partitions.
We can imagine attaching to any molecule a hypothetical string of bits to serve as an address of sorts. One value for the bit string could say that it is within a biological cell, another that it is outside the cell. If more fidelity is desired, one could denote intranuclear, [^14] That is, within the nucleus of a cell. cytosolic, [^15] Within the region between a cell’s nucleus and its outer boundary. or extrecellular. Alternatively (or going further), one could create separate label string values for any given partition of interest, such as one value for the shared space between a group of cells and then additionally one value each for every type of cell as described in the Introduction.
@@
Computational duality allows us to capture richer spatial aspects of evolutionary dynamics in the molecular reformulation of EGT discussed above, as we can account for molecule placement simply by creating multiple entries in $x$ which correspond to the same physical molecular structure across as many partitions as relevant for a given problem. For $q$ cell types of interest, one extracellular compartment, and $n_{0}$ distinct molecular structures of interest, we can have up to $n_{\text{max}}=n_{0}(q+1)$ multicules. It is a set of $n\leq n_{\text{max}}$ multicules which we consider here.
Computational duality is a particularly interesting proposition in that is is capable of converting essentially any collection of disparate types of biological processes into a single coherent process type. Abstracting small-molecule structural information and position as part of a collective state can, in principle, be extended to discretizations of other properties such as conformational states of proteins, methylation states of DNA. Versions of this concept appear in systems theory in various contexts, and biophysics allows all properties to be represented as abstract analogs to molecular structures, and all processes as abstract analogs to chemical reactions. The general nature of this view is the reason for the term multicule rather than particule. This combined state space is of course computationally intractably large if one wished to include the entirety of most given biological processes, but reduced state spaces can always be chosen for the variables most relevant to a given scientific question. Perhaps even more interesting is that from a mathematical perspective generalized multicules provide a single, unified framework for all of biology at the level of molecular and biochemical physiology.
@@
  • Because this study focused on the mathematical and algorithmic aspects of an applied theory problem in mathematical biology, no experimental data were used or produced.
  • No code was created for this study.
  • Likewise, no other resources were generated for this study. The entirety of the work can be found between the paper and the supplemental information.
@@ ACKNOWLEDGMENTS
This work was funded by a research grant from the Stanford University Physician-Scientist Training Program. The study was further supported by grants from the National Institute of Health (17X074, U54CA261717). We would like to thank Mykel Kochenderfer, Arec Jamgochian, and Benjamin van Roy for guidance in literature review and recommendations of relevant papers in the fields of POMDPs and RL, and the former two for helpful feedback on conceptualization in the early stages of the project. We thank Dena Panovska for her thoughtful insights on linking this paper with experimental methods. Additionally, we thank the late Christopher T. Walsh for valuable discussions of underlying concepts in the early stages of the project and for his mentorship. He will be missed.
@@
2: Model-Predictive Control of the Evolutionary Game. This section describes the Bellman operator and Bellman function, particularly as relates to this paper.